Some students simplify:
2x^2 - 32 / x-4 as 2x^2 - 32 / x-4 = 2x-8
Choose values of x to show this is not true. Then simplify it correctly.
May be someone could help me with thise problem.

Answer by stanbon(48545)

Some students simplify:
2x^2 - 32 / x-4 as 2x^2 - 32 / x-4 = 2x-8
Choose values of x to show this is not true.
Let x=0
You get -32/-4 = -8 which is incorrect.
--------------------
Then simplify it correctly.
(2x^2-32)/(x-4) = [2(x-4)(x+4)]/(x-4) = 2(x+4) = 2x+8
==============================
Cheers,
Stan H.